OpenStudy (anonymous):

f(x)=x sqrt(x^2 + 3) + 9 find f'(1) and f''(1). now find the value of x where the 2nd derivative equals zero. (that is, solve the equation f''(x)=0 for x.) x=

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

someones helping me! ahhh!

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4f4fcca5e4b019d0ebaef38c i need help with this too

sam (.sam.):

\[f(x)=x \sqrt{(x^2 + 3)} + 9\] \[f'(x)=(x ^{2}+3)^{1/2}+\frac{x ^{2}}{(x ^{2}+3)^{1/2}}\] \[f'(1)=\frac{5}{2}\]

OpenStudy (anonymous):

what about the other 2?

sam (.sam.):

Differentiate f '(x) again, then equate f ''(x)=0

sam (.sam.):

\[f''(x)=\frac{x}{(x ^{2}+3)^{1/2}}+\frac{x(x ^{2}+6)}{(x ^{2}+3)^{3/2}}\]

sam (.sam.):

f''(x)=0, then solve for x

OpenStudy (anonymous):

can f''(x) be written any other way

OpenStudy (amistre64):

@amistre64 seeing if i can ping myself here

OpenStudy (anonymous):

huh?

OpenStudy (amistre64):

i couldnt find the post, and pinging myself had no effect ...

OpenStudy (amistre64):

@mariomintchev is a ping that sends you a link to where you ping it at

OpenStudy (amistre64):

alot easier to follow to then a link in a fan testy

OpenStudy (amistre64):

writting f'' in another way doesnt change the math; so going thru all the extra work to "simplify" is usually a lost cause

OpenStudy (anonymous):

my site isnt accepting my answer

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

i am lost. which part is needed? second deriviative?

OpenStudy (anonymous):

yes. everything else i have.

OpenStudy (anonymous):

well i need the first derivative first

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

\[\frac{2x^2+3}{\sqrt{x^2+3}}\] if i am not mistaken

OpenStudy (anonymous):

i mean that is the first derivative. second one in a minute

OpenStudy (anonymous):

keep going

OpenStudy (anonymous):

\[\frac{x(2x^2+9)}{(x^2+3)^{\frac{3}{2}}}\]

OpenStudy (anonymous):

now this will only by zero if x = 0

OpenStudy (anonymous):

since \[2x^2+9\geq 9\] so that part is never zero

OpenStudy (anonymous):

that should do it right?

OpenStudy (anonymous):

it keeps saying its wrong. maybe im putting it in wrong. idk but im going nuts.

OpenStudy (amistre64):

use the wolf ;)

OpenStudy (amistre64):

the dirigible is right

OpenStudy (anonymous):

its not accepting that answer

OpenStudy (amistre64):

make sure youve given us the correct f(x)

OpenStudy (amistre64):

retype it for us

OpenStudy (anonymous):

what SAM has above is correct. that is what the problem looks like

OpenStudy (amistre64):

and you say it wasnts f'(1) and f''(1) right?

OpenStudy (anonymous):

scroll up ^^^^

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for f'(1) its 5/2.

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

f''(1) = 11/8

OpenStudy (anonymous):

o my god! yes! please show me how you got that

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=2nd+derivative+%28x*sqrt%28x%5E2+%2B+3%29+%2B+9%29%2C+x%3D1 you just plug in x=1 into f''

OpenStudy (amistre64):

\[\frac{1(2(1)^2+9)}{((1)^2+3)^{\frac{3}{2}}}\] \[\frac{11 }{(4)^{\frac{3}{2}}}\] \[\frac{11 }{(64)^{\frac{1}{2}}}\] \[\frac{11 }{8}\]

OpenStudy (anonymous):

ok ur awesome! need a final answer for 2 more problems like this so if u could give me a hand that would be sweet!

OpenStudy (anonymous):

i already posted them long ago. i just dont have a final answer.